OpenStudy (anonymous):

n the function f(x) = 5x - 11 how does a decrease of 1 in the independent variable x affect f(x)?

jimthompson5910 (jim_thompson5910):

say x = 1 f(x) = 5x - 11 f(1) = 5(1) - 11 f(1) = 5 - 11 f(1) = -6 So when x = 1, f(x) = -6

OpenStudy (anonymous):

f(x) will have a decrease of 5 f(x) will have an increase of 5 f(x) will have an increase of 16 f(x) will have a decrease of 5x - 5 are the choices

jimthompson5910 (jim_thompson5910):

now say x = 0 (x decreases by 1) f(x) = 5x - 11 f(0) = 5(0) - 11 f(0) = 0 - 11 f(0) = -11 Now when x = 0, f(x) = -11

jimthompson5910 (jim_thompson5910):

you have these two facts x = 1, f(x) = -6 and x = 0, f(x) = -11

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

what does that have to do with the problem

jimthompson5910 (jim_thompson5910):

going from f(x) = -6 to f(x) = -11 has f(x) decreasing by _____

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

so f(x) decreases by 5 when x decreases by 1

jimthompson5910 (jim_thompson5910):

which is exactly what the slope is saying

OpenStudy (anonymous):

is f(x) will have a decrease of 5 the answer

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

you sure

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks very much